density of vegetable oil with density 921 kg/m3 is found to have a mass of 34.9g. Find the volume of this sample
1
Expert's answer
2018-09-03T15:41:08-0400
First you need to transfer the mass of the sample from grams to kilograms (1 kilogram is 1000 grams):
34.9 / 1000 = 0.0349 kg.
The formula for finding the density of the sample is as follows:
ρ=m/V
m – a mass of the sample, V – a volume, ρ – a density.
Based on this formula, the sample volume is as follows:
V= m/ρ= 0.0349/921=0.0000379 m^3=37.9 〖cm〗^3